OpenStudy (anonymous):

Prove, in complete sentences, whether the left expression is equal to the right expression and discuss which property applies. 2(x − 3) 2x − 6 HELP! ILL GIVE A MEDAL

OpenStudy (anonymous):

"2(x − 3) = 2x − 6" 1. Apply the distributive property. 2x - 6 2 times x = 2x 2 times -3 = -6 2. 2x - 6 = 2x − 6 The left expression is equal to the right expression. The distributive property applies

OpenStudy (anonymous):

So Is That The Answer? ;o

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Thanks :D

OpenStudy (anonymous):

no problem